Write an equation of the line in slope-intercept form that is parallel to the given line and passes through the given point
y=5x + 11 (8,0)

slope______
equation_________
Thanks

Found 2 solutions by stanbon, algebrahelp101:
Answer by stanbon(75887)

The slope is 5.
Since it is a line it has the form y=mx+b
Replace m, x, and y with m=5, x=8, y=0 to solve for "b"
0=5(8)+b
b=-40
The equation you want is y=5x-40
Cheers
Stan H.

Answer by algebrahelp101(11)

The slope is 5.
Since it is a line it has the form y=mx+b
Replace m, x, and y with m=5, x=8, y=0 to solve for "b"
0=5(8)+b
b=-40
The equation you want is y=5x-40
Cheers